# SynLlama: Generating Synthesizable Molecules and Their Analogs with Large Language Models

### v1

@@ 2.4 SynLlama Model Benchmarks
Since we are formulating the synthetic tasks using purely language-based modeling, where all reactions are expressed in SMARTS templates and molecules in SMILES strings, it is extremely important to quantify the capacity of SynLlama models for instruction following and comprehension of reaction chemistry. To assess SynLlama’s capability of following instructions, we select three benchmarking criteria as shown in Table 1. The first one is “Valid JSON,” which examines whether the output format is a parsable JSON following the fine-tuned templates that will be necessary for the downstream reconstruction algorithm. The second criterion is “Template Memorization,” which assesses the model’s ability to memorize the 91 reaction templates we use to define our custom synthesizable chemical space. Lastly, we benchmark on “BB Selection,” which evaluates whether the “building blocks” section in the responses can accurately identify and select all the building blocks, namely reactants across every reaction in a single response that are not a product of another reaction from the ”reactions” section of the responses.
@@
To assess SynLlama’s comprehension of reaction chemistry, we focus on individual reactions and summarize the three critical aspects as: (1) the percentage of “Valid SMILES” out of all SMILES strings in the responses, which is essential for assessing SynLlama’s learning outcome of string-based chemical representations in general, (2) the percentage of “Matched Reactants”, which calculates whether the generated reactants match the reactant templates specified in the predicted reactions, and (3) the percentage of “Good Products,” which assess if the predicted product can indeed be generated by applying the proposed reaction templates onto the reactants. Overall, these three benchmarks can collectively assess SynLlama’s capability to perform chemical reactions in string representations.
@@
In Table 1, all four trained SynLlama models are evaluated on both in-distribution training data and out-of-distribution testing and ChEMBL 19 data to assess the benchmarks outlined above. In the instruction-following benchmarks, each model exhibits strong adherence (over 95 percent) to the fine-tuned response structure across all datasets. This impressive performance indicates that fine-tuning effectively retains the specified output structure when trained with over 100,000 samples. Furthermore, all four models successfully memorize the 91 provided RXN templates and select the building blocks (BBs) from all predicted reactants over 99 percent of the time. This capability further enhances the coupling effectiveness of the downstream reconstruction algorithm with the SynLlama raw output, as it only requires information about reaction sequences and predicted building blocks.
@@
In the reaction chemistry benchmarking results, a clearer trend emerges: models, irrespective of their parameter size, show improved comprehension of reaction chemistry on all three datasets as the amount of training data increases. Notably, all four models maintain their performance from training to testing data but exhibit a decline in “Matched Reactants” and “Good Products” performance when generalizing to the ChEMBL data. The reason behind this is that the testing data is generated in the same manner as the training data but with a different set of building blocks, whereas the ChEMBL dataset occupies a distinctly different chemical space, as previously noted by Gao et al. 48. Despite reductions in their performance for ChEMBL molecules as shown in Supplementary Figure S2, SynLlama-8B-500k and SynLlama-1B-2M can still generate complete and valid syntheses over 50% of the time. This indicates that SynLlama’s raw results alone have potential utility for synthesis planning, even without additional processing through the reconstruction algorithm.
@@
Testing Data Valid JSON 95.90% 97.50% 96.60% 97.30%
Template Mem. 99.97% 100.0% 99.95% 100.0%
BB Selection 99.68% 99.98% 99.79% 99.91%
Valid SMILES 94.31% 99.63% 94.81% 99.66%
Matched Reactants 77.13% 96.39% 81.82% 97.89%
Good Products 78.97% 96.44% 80.59% 97.92%
@@
Table 1: Benchmarks of SynLlama inferences on 1000 training, testing, and ChEMBL data. We first select 1000 SMILES strings from the training examples, testing examples and the ChEMBL dataset, and then run inferences using SynLlama models. The detailed descriptions of each benchmark can be found in the main text. Here, we run SynLlama inferences at $T=0.1$ and $TopP=0.1$ to generate reproducible benchmarking results (see Supplementary Table S1).
When comparing SynLlama-8B-500k and SynLlama-1B-500k, we observe that the larger model demonstrates better performance when trained on the same amount of data. While additional training data could further improve the 8B model based on the current trend, its higher computational cost makes this pursuit less practical. However, as the fine-tuning computational costs for for SynLlama-8B-500k and SynLlama-1B-2M require approximately the same A40-GPU hours, and given the comparable benchmark performance between them, we decided to move forward with SynLlama-1B-2M, simplified as SynLlama, for the subsequent tasks due to its faster inference speed.
@@ 3.1 Synthesis Planning for Unseen Molecules
Having demonstrated that SynLlama models can reliably predict reaction sequences and building blocks, we now employ SynLlama to plan the synthesis of previously unseen molecules from the Testing and ChEMBL data as well as a set of Enamine compounds. Initially, we follow the standard reconstruction approach within the Enamine chemical space: SynLlama generates raw output predictions in a retrosynthetic manner and this output then gets processed through the reconstruction algorithm to create target input molecules or analogs using building blocks exclusively from the Enamine library. This evaluation method aligns with that used by existing approaches such as SynNet48, ChemProjector49, and Synformer50, that we compare with in regards the number of reconstructions of the target molecule from Enamine BBs in Table 2. In this comparison SynNet48 only serves as a baseline for coverage of synthesizable chemical space as already discussed elsewhere.49 Instead we compare SynLlama against more recent algorithms such as ChemProjector49 and Synformer50 that use more powerful architectures and much larger training data to achieve substantially better performance than SynNet. Furthermore Synformer also benefits from an expanded set of 115 reaction templates that includes tri-molecular reactions compared to the 91 unimolecular and bi-molecular reactions used by SynNet, ChemProjector, and SynLlama.
Table 2 shows that the SynLlama model greatly outperforms ChemProjector49 in target reconstruction with Enamine BBs on all three data sets, and at the same time reduces data requirements by over 60-fold. In fact Supplementary Table S2 shows that SynLlama directly predicts more target molecule synthesis without any need to project into the Enamine BB space compared to ChemProjector. SynLlama also performs competitively against Synformer in the Enamine synthesis space while reducing data requirements by over 40-fold, and thus the gains made by Synformer come primarily from the use of ternary reaction templates not yet implemented in SynLlama.
In addition to reducing large data requirements for training, an additional benefit of SynLlama is to reconstruct target molecules with new types of commercially available predicted BBs, i.e. beyond the Enamine BBs, even though it was not explicitly trained for this purpose. Hence new synthesis pathways involving New BBs can increase the reaction pathways for the same or additional target molecules in all three data sets as seen in Table 2. Because some target molecules can be synthesized by either using only Enamine BBs or with additions of New BBs, the Total number of reconstructions refers to only the unique reaction targets synthesized which increases SynLlama reconstruction rate to 64.2%, 55.8%, and 21.6% for the 1000 molecules in each of the Testing, Enamine Set, and ChEMBL data sets, respectively. It is especially noteworthy that the New BBs greatly improves the synthesizability of the more difficult ChEMBL data set in particular, such that it slightly exceeds the performance of Synformer in spite of the fact that SynLlama has fewer reaction templates.
Finally, in some cases the target molecule can not be reconstructed, or a synthesis pathway for a related analog molecule is the actual desired use case. Hence we assess analog quality using three molecular similarity scores between the target molecule and its most similar analog: (1) Tanimoto similarity based on 4096-bit Morgan fingerprints of both molecules, (2) 4096-bit Morgan fingerprints of Murcko scaffolds, and (3) Gobbi 2D pharmacophore fingerprints of both molecules. Full reconstruction of a target molecule results in a similarity score of 1 across all metrics. Table 2 and Supplementary Table S3 shows that SynLlama is superior to ChemProjector in generating synthetic pathways to more similar analogs to the target molecule. However compared to Synformer the SynLlama model shows degradation in analog similarity that is likely an issue of having $\sim$25fewer reaction templates.
Dataset Method # of Recon. Mol. Similarity
Enamine BB New BB Total Morgan Scaffold Gobbi
Testing Data SynNet48 107 - 107 0.46 0.51 0.35
ChemProj.49 304 - 304 0.78 0.77 0.85
SynLlama 616 123 642 0.94 0.94 0.95
Enamine Set 1 SynNet48 110 - 110 0.52 - -
ChemProj.49 462 - 462 0.82 0.85 0.83
Synformer50 660 - 660 0.91 0.92 0.89
SynLlama 498 61 558 0.87 0.88 0.85
ChEMBL Data SynNet 54 - 54 0.43 0.20 0.27
ChemProj. 133 - 133 0.60 0.59 0.56
Synformer50 198 - 198 0.67 0.72 0.72
SynLlama 160 84 216 0.66 0.67 0.62
@@
Table 2: Comparison of synthesis planning performance among methods. The Testing, Enamine Set 1, and ChEMBL data sets are comprised of 1000 unseen molecules each. SynNet is a baseline performer as discussed by Luo and co-workers.49 Synformer is excluded from the Testing set comparison because it trains on all available BBs without any splitting to separate out testing BBs. The number of total training data used for: ChemProjector (128M), Synformer (85M), and SynLlama (2M). SynLlama is trained with greedy hyperparameters as described in Supplementary Table S1. We use Molport64 to determine whether SynLlama-predicted BBs are purchasable, with more detailed information reported in the Supplementary Information. The similarity scores include all analog molecules with successful synthesis pathways, as well as successfully reconstructed target molecules. See Supplementary Table S3 for evaluated similarity metrics for only analog molecules.
@@ 3.2 Synthesizable Analog Search for De Novo Molecules
Previous research has highlighted the fact that molecules proposed by generative models struggle in their practical synthesizability39, 40, 41, even when the proposed molecules are optimized by empirical heuristics of a molecule’s synthetic accessibility (SA)42. In this section we illustrate SynLlama’s potential to bridge the gap between generative molecule design and practical synthesis, in which SynLlama is coupled with molecular candidates from a structure-based generative model, iMiner30 performed on the SARS-CoV-2 Main Protease (Mpro). The iMiner algorithm searches through chemical space for drug molecules by optimizing the molecular docking score and drug likeliness, but does not guarantee any level of synthesizability of new proposed molecules; further details in regards iMiner are provided in the Supplementary Information.
@@
Figure 2: SynLlama performance on generating synthesizable analogs for iMiner proposed binders to SARS-CoV2 Mpro. (a). Correlation plot comparing docking scores of original iMiner molecules and the average docking scores of ten most similar analogs (based on Morgan Fingerprints) proposed by SynLlama. Each data point is color-coded by the average Morgan fingerprint similarity computed between iMiner molecules and analogs, with brighter being less similar and darker being more similar. The shaded area represents an energy uncertainty range of ±$\pm 2kcal/mol$, which is the typical uncerntainty for AutoDock Vina scores. (b). SA score distribution of iMiner molecules and SynLlama-proposed analogs. (c). Example iMiner molecule and one of its analog proposed by SynLlama. Using the iMiner molecule as input, SynLlama successfully proposes an analog that retains docking score with decreased SA score. For each binder, we average scores from the top 10 structurally similar molecules for a meaningful distributional analysis. (d) Synthesis pathway of the analog molecule of an iMiner de novo molecule in (c) using SynLlama BBs via indole synthesis (RXN 23) and amide coupling (RXN 76) reaction templates.
@@
Here we select the top 500 iMiner-generated binders based on their AutoDock Vina65 docking scores to Mpro (PDB: 7LTJ)66), and process them through SynLlama to generate synthesizable analogs constrained to Enamine BBs, for which we also perform docking to obtain their best docking scores. In Figure 2(a), the RMSE of docking scores of the generated analogs with respect to the original iMiner molecules is 1.04 kcal/mol, which is an acceptable range of inherent docking score errors reported by Trott and co-workers65. Furthermore, as demonstrated in Figure 2(b), the SA score distribution of the generated analogs show a notable decrease from de novo generated binders, demonstrating improved synthetic accessibility of the analogs proposed by SynLlama. Figure 2(c) shows an example of an iMiner molecule and one of its analog proposed by SynLlama, and Figure 2(d) demonstrates the needed BBs and reaction templates to synthesize the napthalene analog using SynLlama.
@@ 3.3 Local Hit Expansion for Binder Molecules
Because SynLlama breaks down the original target molecule for synthesis into building blocks, by nature this method allows for diverse exploration around parts of the molecular scaffold rather than only on a whole target molecule. In a final task, we apply SynLlama to expand on a hit molecule for the same SARS-CoV-2 Mpro protein (PDB: 7LTJ)66 to discover synthesizable molecules that have better relative binding free energies (RBFEs) confirmed by both experiments and accurate free energy perturbation (FEP) calculations.
Figure 3: Hit expansion of binders to SARS-CoV-2 Mpro (PDB: 7LTJ) with SynLlama. (a) Target hit molecule and the predicted synthetic pathway to generate analogs of the target binder proposed by SynLlama. The core scaffold of the hit molecules are colored in red. (b) FEP benchmarking on the protein system 7LTJ. The experimental IC50 values are converted to the binding free energies, $\Delta G$, which shows good correlations with the $\Delta G$ calculated by FEP. (c) Hit compound and proposed analogs from SynLlama. The red-colored fragment represents the hit compound. The blue-colored fragment represents one SynLlama-generated analog that was also experimentally verified to have better binding affinity than the hit molecule. The green-colored fragments represent six potential binders that are within the 1 kcal/mol uncertainty range of the FEP score of the original hit.
As shown in Figure 3(a), the hit molecule has a core scaffold of uracil and ortho-dichlorobenzene connected by a piperizine linker. Inspired by an experimental hit expansion campaign by Kneller et al.67, we follow their practices to propose only functional group substitutions on the benzene ring while keeping the linker and uracil intact. Specifically, we use SynLlama to generate 50 synthesizable analogs constrained to only Enamine BBs of the hit compound and filter for molecules that only modifies the benzene ring, harvesting a total of 8 analog molecules that fulfill the criteria. The analogs are placed in a pose configuration similar to the original hit molecule and we perform FEP on all 8 molecules. As shown in Figure 3(b), all of the calculated FEP values correlate well with experimental $\Delta G$ converted from IC50 and are all within ±$\pm 1kcal/mol$ energy error of experimental binding affinities. Then as shown Figure 3(c), SynLlama rediscovers two molecules from the reported hit expansion campaign screening, both of which have experimentally validated IC50 values; in fact one molecule highlighted in blue was identified as the best compound from the hit-expansion efforts reported by Kneller et al.67. The six molecules highlighted in green also demonstrate improved FEP-calculated binding affinities compared with the lead molecule, suggesting new directions for hit optimization on this ligand series.

### v2

@@ 2.4 SynLlama Model Benchmarks
Since we are formulating the synthetic tasks using purely language-based modeling, where all reactions are expressed in SMARTS templates and molecules in SMILES strings, it is extremely important to quantify the capacity of SynLlama models for instruction following and comprehension of reaction chemistry. To assess SynLlama’s ability to follow instructions, we select three benchmarking criteria as shown in Table 1. The first is “Valid JSON,” which examines whether the output format is a parsable JSON following the fine-tuned templates that will be necessary for the downstream reconstruction algorithm. The second criterion is “Template Memorization,” which assesses the model’s ability to memorize the provided reaction templates that define our synthesizable chemical space. Lastly, we benchmark on “BB Selection,” which evaluates whether the “building blocks” section in the responses can accurately identify and select all the building blocks from the “reactions” section of the responses.
@@
To assess SynLlama’s comprehension of reaction chemistry, we focus on individual reactions and summarize the three critical aspects as: (1) the percentage of “Valid SMILES” out of all SMILES strings in the responses, which is essential for assessing SynLlama’s learning outcome of string-based chemical representations in general, (2) the percentage of “Matched Reactants,” which calculates whether the generated reactants match the reactant templates specified in the predicted reactions, and (3) the percentage of “Good Products,” which assess if the predicted product can indeed be generated by applying the proposed reaction templates onto the reactants. Overall, these six benchmarks can collectively assess SynLlama’s capability to follow instructions and perform chemical reactions in string representations.
@@
In Table 1, all four trained SynLlama models are evaluated on both in-distribution training data and out-of-distribution testing and ChEMBL 19 data to assess the benchmarks outlined above. In the instruction-following benchmarks, most models exhibit strong adherence (over 90 percent) to the fine-tuned response structure across all datasets. This impressive performance indicates that fine-tuning effectively retains the specified output structure when trained with over 100,000 samples. Furthermore, all four models successfully memorized the provided RXN templates and selected the building blocks (BBs) from all predicted reactants over 99 percent of the time. This capability further enhances the coupling effectiveness of the downstream reconstruction algorithm with the SynLlama raw output, as it only requires information about reaction sequences and predicted building blocks.
@@
In the reaction chemistry benchmarking results, a clearer trend emerges: models, regardless of their parameter size, show improved comprehension of reaction chemistry in all three datasets as the amount of training data increases. Notably, most models maintain their performance from training to testing data, but exhibit a greater decline in “Matched Reactants” and “Good Products” performance when generalizing to the ChEMBL data. The reason behind this is that the testing data are generated in the same manner as the training data but with a different set of building blocks, while the ChEMBL data occupies a different chemical space, as previously noted by Gao et al. 48. Despite the reductions in their performance for ChEMBL molecules, as shown in Supplementary Figure S2, SynLlama-8B-500k and SynLlama-1B-2M can still generate complete and valid syntheses over 50% of the time without any downstream processing. These results indicates that SynLlama’s raw results alone have potential utility for synthesis planning for unseen drug-like molecules.
@@
Testing Data Valid JSON 91.00% 94.20% 88.70% 93.90%
Template Mem. 99.91% 100.0% 99.97% 100.0%
BB Selection 99.75% 99.96% 99.73% 99.98%
Valid SMILES 94.37% 99.13% 87.66% 99.50%
Matched Reactants 77.51% 94.83% 65.63% 96.90%
Good Products 74.11% 94.16% 69.54% 96.39%
@@
Table 1: Benchmarks of SynLlama inferences on 1000 training, testing, and ChEMBL data. We first select 1000 SMILES strings from the training examples, testing examples and the ChEMBL dataset, and then run inferences using SynLlama models trained on RXN Set 1. Benchmarking result for SynLlama models trained on RXN Set 2 can be found in Table S2. The detailed descriptions of each benchmark can be found in the main text. Here, we run SynLlama inferences at $T=0.1$ and $TopP=0.1$ to generate reproducible benchmarking results (see Supplementary Table S1).
When comparing SynLlama-8B-500k and SynLlama-1B-500k, we observe that the larger model demonstrates better performance when trained on the same amount of data. Although additional training data could further improve the 8B model based on the current trend, its higher computational cost makes this pursuit less practical. However, as the fine-tuning computational costs for SynLlama-8B-500k and SynLlama-1B-2M require approximately the same A40-GPU hours, and given the comparable benchmark performance between them, we decided to move forward with SynLlama-1B-2M, simplified as SynLlama, for the subsequent tasks due to its faster inference speed.
@@ 3.1 Synthesis Planning for Unseen Molecules
Dataset Method # of Recon. Mol. Morgan Sim.
Enamine BB New BB Total
Testing-Set 1 SynNet48 107 - 107 0.46
ChemProjector49 301 - 301 0.79
SynLlama(RXN Set 1) 631 125 669 0.93
Testing-Set 2 SynLlama(RXN Set 2) 538 114 582 0.91
Enamine Data SynNet 110 - 110 0.57
ChemProjector 462 - 462 0.82
Synformer50 660 - 660 0.91
SynLlama(RXN Set 1) 527 100 568 0.87
SynLlama(RXN Set 2) 691 232 741 0.92
ChEMBL Data SynNet 54 - 54 0.43
ChemProjector 133 - 133 0.60
Synformer 198 - 198 0.67
SynLlama(RXN Set 1) 165 95 223 0.66
SynLlama(RXN Set 2) 197 152 287 0.68
@@
Table 2: Comparison of synthesis planning performance among different methods. Both Testing, Enamine, and ChEMBL data sets are comprised of 1000 unseen molecules each. Details of each benchmark are described in the main text. The Morgan similarity scores include all analog molecules with successful synthesis pathways, as well as successfully reconstructed target molecules. The number of total training data and reaction set each method used for is: SynNet (200K, RXN Set 1) ChemProjector (128M, RXN Set 1), Synformer (85M, RXN Set 2), and SynLlama (2M, RXN Set 1 & 2). Inference details and identification of purchasable New BBs with Molport63 are described in the Supplementary Information.
Having demonstrated that SynLlama models can reliably predict reaction sequences and building blocks, we now use SynLlama to plan the synthesis of two groups of previously unseen molecules. The first group comprises two sets of 1000 testing molecules, generated similarly to the training data using testing building blocks and corresponding reaction templates. The second group includes two sets of 1000 molecules extracted from the publicly available ChEMBL database19 and the Enamine diversity set20.
Initially, we follow the standard reconstruction approach within the Enamine chemical space: SynLlama generates raw output predictions in a retrosynthetic manner, and this output then gets processed through the reconstruction algorithm to create target input molecules or analogs using building blocks exclusively from the Enamine library. This evaluation method aligns with that used by existing approaches such as SynNet48, ChemProjector49, and Synformer50, and we compare these methods on the number of reconstructions of the target molecule with Enamine BBs in Table 2. In this comparison, SynNet48 serves as a baseline for synthesizable chemical space coverage.49 Furthermore, when trained on their respective reaction sets and using just Enamine BBs along, SynLlama outperforms ChemProjector49 and matches the performance of Synformer50, while reducing the number of training data by 60-fold and 40-fold.
In addition to reducing data requirements for training, SynLlama can also reconstruct target molecules using commercially available BBs beyond Enamine, even without specific training for this purpose. These ‘New BBs’ add possible synthetic pathways to reconstruct target molecules in all four datasets, as seen in Table 2 and detailed in Tables S3 and S4. Since some target molecules can be synthesized through multiple pathways, either using only Enamine BBs or with the addition of New BBs, the ‘Total’ column in the table reflects the number of unique target molecules reconstructed with SynLlama. With these New BBs, SynLlama’s best reconstruction rates increase to 66.9%, 58.2%, 74.1%, and 28.7% for the 1000 molecules in both Testing Sets, Enamine, and ChEMBL data, respectively. These results clearly outperform previous methods that lack such search capacity of novel building blocks. It is especially noteworthy that New BBs significantly enhance synthetic accessibility of unseen molecules, suggesting that SynLlama learns reaction chemistry well enough to predict novel building blocks.
Finally, when the target molecule cannot be reconstructed, we assess the quality of the analog using three molecular similarity scores between the target molecule and its most similar analog: (1) Tanimoto similarity based on 4096-bit Morgan fingerprints64, (2) 4096-bit Morgan fingerprints of Murcko scaffolds65, and (3) Gobbi 2D pharmacophore fingerprints66. In table 2, the similarity metrics reported are average values of all trial molecules, including those that are fully reconstructed. In 2, we only show the first similarity metric due to the space limit. More details on the metrics and results for molecules that cannot be reconstructed are explored in Tables S5 and S6. Overall, these results collectively show that SynLlama is capable of planning synthesis for unseen molecules or their analogs with purchasable building blocks, highlighting its potential use in drug discovery.
@@ 3.2 Synthesizable Analog Search for De Novo Molecules
Previous research has shown that molecules proposed by generative models often face challenges in practical synthesizability39, 40, 41, even when optimized using empirical heuristics for synthetic accessibility. In this section, we demonstrate SynLlama’s potential to bridge the gap between generative molecule design and practical synthesis. Specifically, SynLlama is integrated with candidates from a structure-based generative model, iMiner30, which optimizes drug-likeness and molecular docking scores against the SARS-CoV-2 Main Protease (Mpro). More details in regards to iMiner are provided in the Supplementary Information.
@@
Figure 2: SynLlama performance on generating synthesizable analogs for iMiner proposed binders to SARS-CoV2 Mpro. (a). Correlation plot comparing docking scores of iMiner molecules and the average docking scores of ten most similar analogs (based on Morgan Fingerprints) proposed by SynLlama trained on RXN Set 2. Each data point is color-coded by the average Morgan fingerprint similarity computed between iMiner molecules and analogs, with brighter being more similar and darker being less similar. The shaded area represents an energy uncertainty range of ±$\pm 2kcal/mol$, which is the typical uncerntainty for AutoDock Vina scores. (b). SA score distribution of iMiner molecules and SynLlama-proposed analogs. (c). Example iMiner molecule and one of its analog proposed by SynLlama. Using the iMiner molecule as input, SynLlama successfully proposes an analog that retains docking score with decreased SA score. For each binder, we average scores from the top 10 structurally similar molecules for a meaningful distributional analysis. (d) Synthesis pathway of the analog molecule of an iMiner de novo molecule in (c) using SynLlama BBs via indole synthesis (RXN 23) and amide coupling (RXN 76) reaction templates.
Here, we select the top 500 iMiner-generated binders based on their AutoDock Vina67 docking scores to Mpro (PDB: 7L1168), and process them through SynLlama, trained on RXN Set 2, to generate synthesizable analogs constrained to Enamine BBs. For the top-10 generated analogs, we perform molecular docking with the same procedure and report their average docking score and average Morgan fingerprint similarity in Figure2. We choose the SynLlama model trained on RXN Set 2 for their superior performance in analog generation explored in the previous section. A similar distribution analysis of the iMiner analogs generated with SynLlama trained on RXN Set 1 is presented in Figure S3.
@@
In Figure 2(a), the RMSE of docking scores of the generated analogs with respect to the original iMiner molecules is 1.04 kcal/mol, which is an acceptable range of inherent docking score errors reported by Trott et al67. Furthermore, as demonstrated in Figure 2(b), the SA score distribution of the generated analogs show a notable decrease from de novo generated binders, demonstrating improved synthetic accessibility of the analogs proposed by SynLlama. These results suggest that SynLlama can generate synthetic pathways for iMiner-proposed molecules or their analogs, enhancing synthetic accessibility of de novo molecules when combined with generative models.
@@ 3.3 Local Hit Expansion for Binder Molecules
Because SynLlama breaks down the original target molecule for synthesis into building blocks, by nature this method allows diverse exploration around parts of the molecular scaffold rather than only on a whole target molecule. In a final task, we apply SynLlama to expand on a hit molecule for the same SARS-CoV-2 Mpro protein (PDB: 7LTJ)69 to discover synthesizable molecules that have better relative binding free energies (RBFEs) confirmed by both experiments and accurate free energy perturbation (FEP) calculations.
Figure 3: Hit expansion of binders to SARS-CoV-2 Mpro (PDB: 7LTJ) with SynLlama. (a). Target hit molecule and the predicted synthetic pathway to generate analogs of the target binder proposed by SynLlama. The core scaffold of the hit molecules are colored in red. (b). Hit compound and proposed analogs from SynLlama. The red-colored fragment represents the hit compound. The blue-colored fragment represents one SynLlama-generated analog that was also experimentally verified to have better binding affinity than the hit molecule. The green-colored fragments represent six potential binders that are within the 1 kcal/mol uncertainty range of the FEP score of the original hit.
As shown in Figure 3(a), the hit molecule has a core scaffold of uracil and ortho-dichlorobenzene connected by a piperizine linker. Inspired by an experimental hit expansion campaign by Kneller et al. 70, we follow their practices to propose only functional group substitutions on the benzene ring while keeping the linker and uracil intact. Specifically, we use SynLlama to generate 50 synthesizable analogs constrained to only Enamine BBs of the hit compound and filter for molecules that only modifies the benzene ring, harvesting a total of 8 analog molecules that fulfill the criteria. The analogs are placed in a pose configuration similar to the original hit molecule and we perform FEP on all 8 molecules.
To verify the FEP results, we first choose around 10 synthesized and experimentally tested molecules to benchmark the accuracy of FEP for this system. In Figure S4, all calculated FEP values show a good correlation with experimental $\Delta G$ converted from IC50, with an average RMSE of less than $1kcal/mol$. After this validation, we run FEP for the eight proposed molecules to assess their binding affinities. As shown Figure 3(b), SynLlama rediscovers two molecules from the reported hit expansion campaign, both with experimentally validated IC50 values. Notably, one molecule highlighted in blue was identified as the best compound from the expansion efforts reported by Kneller et al.70. The other six molecules highlighted in green also demonstrate desirable FEP-calculated binding affinities compared to the lead molecule, suggesting new directions for hit optimization in this ligand series. These results show that SynLlama effectively explores local chemistry with purchasable building blocks, saving time and effort for medicinal chemists who would otherwise manually enumerate and check availability at different steps.